OpenStudy (anonymous):

One who and one what cost 12 dollars. One who and one where cost 11 dollars. One what and one where cost ten dollars. How many dollars do 2 who’s cost?

RhondaSommer (rhondasommer):

who=x what=y where=a x+y=12 x+a=11 y+a=10 add the first 2 together so it is 2 who's and a what and a where subtract the last equation so it only leaves 2 who's :)